# Efficient Architectures for High Resolution Vision-Language Models

### v1

@@ 1 Introduction
The integration of visual capabilities as extensions to Large Language Models (LLMs) has led to the emergence of large Vision-Language Models (VLMs), currently excelling in tasks like image captioning and visual question answering. Notable examples include Flamingo (Alayrac et al., 2022) or Monkey (Li et al., 2023d), in this last case bringing improvements to tasks that involve scene-text by processing higher-resolution images. LLaVA-NeXT (Liu et al., 2024) followed Monkey’s lead with similar enhancements, but at the cost of a quadratic increase in the computational complexity of the language model, as it skipped a resampler module that compresses large sequence lengths, from the encoder, into a fixed size of query vectors.
Recent focus has shifted towards smaller VLMs that can run on hardware-constrained devices. Models like MoE-LLaVA (Lin et al., 2024) or moondream2 (Vikhyat, 2024) achieve impressive performance with a fraction of the parameters of their predecessors. This work further advances small VLMs with Pheye, i.e. a family of compact models that can process high-resolution images with fewer parameters and computational demands, expanding VLM applications to resource-limited environments where understanding fine details is crucial.
Specifically, Pheye employs a frozen instruction-tuned language model (Li et al., 2023c) in conjunction with a frozen pre-trained CLIP (Radford et al., 2021) vision model, which are linked by dense cross-attention layers inserted before the language model’s layers. To process high-resolution images, we use two sets of LoRA (Hu et al., 2021) adapters in the vision encoder, one for the global image and another for local high-resolution patches.
@@
Notably, Pheye is competitive with similarly sized models, particularly in tasks involving scene-text, such as TextVQA (Singh et al., 2019). Pheye requires only a fraction of the training parameters, being more efficient in connecting the vision and language modalities, and in processing input images at higher resolutions. The source code associated to the experiments reported on this paper, as well as the trained models, are available from a public GitHub repository [^1] https://github.com/miguelscarv/pheye.
@@ 2.1 The Proposed Architecture
The vision encoder consists of a ViT with two sets of LoRA adapters, one for encoding global images and another for local high-resolution patches, as outlined in Appendix B.2. To better distinguish between global and local patch embeddings and improve convergence, we introduce two LayerNorm (Ba et al., 2016) layers after the ViT, and before adding learned positional embeddings. These layers are applied separately, respectively processing the global and local patch embeddings.
@@
Given the computational efficiency, strong task performance, and reduced number of trainable parameters, we use cross-attention to combine modalities while keeping the language model frozen. Inspired by the Flamingo architecture (Alayrac et al., 2022), we replace vanilla cross-attention modules with dense cross-attention modules, inserting them at regular intervals before the decoder layers, as shown in Figure 2. This design was motivated Flamingo’s demonstration that gated cross-attention outperforms vanilla cross-attention even when parameter counts are equal. However, we deviate from Flamingo’s gated cross-attention layers, as early experiments showed that the gating mechanism hindered convergence. To preserve the language model’s initial integrity, we initialize the output matrices of the dense cross-attention modules with values sampled from a normal distribution with a mean of zero and a variance approaching zero. This way, the cross-attention layers have a minor influence during the initial training epochs.
@@ 2.2 Analysis of the Computational Complexity
$\vspace{-0.01cm}\mathrm{\mathbb{T}_{ViT}}=4ND^{2}+DN^{2}+8ND^{2}.\vspace{-0.05cm}$ (1)
@@
We can compare the efficiency of our vision encoder against a ViT that operates at a resolution of 672$\times$672 pixels with a patch size of 14 pixels, resulting in $N=2305$ input tokens. Our method would process 10 images, where 1 is global and 9 are local sub-images, at a resolution of 224$\times$224 pixels with a patch size of 14, yielding $N^{\prime}=2570$ tokens. With $D=1280$, our method is approximately 1.02 as efficient as the alternative. While this corresponds to little to no improvement, our method does offer a significant advantage: it eliminates the need to fine-tune the underlying ViT, allowing us to focus solely on training the LoRA parameters to increase the input resolution.
@@ 3.1 Experimental Setup
Our experiments aimed to assess the effects of increasing the input image resolution, and to examine the impact of augmenting the frequency of dense cross-attention layers in the language model. To do this, we compare four model settings, varying the image input resolution between 448$\times$448 pixels and 672$\times$672 pixels, and adjusting the dense cross-attention interval to every 4 or 2 decoder layers.
In all four settings, we used a Phi 1.5 language model (Li et al., 2023c), finetunned on the SlimOrca (Lian et al., 2023) instruction dataset. The ViT is initialized from a CLIP-ViT-H-14 model finetunned on DFN-5B (Fang et al., 2023).
The models were trained in separate stages, similarly to MoE-LLaVA (Lin et al., 2024), using a cross-entropy loss over the output tokens. The different stages are described next, while Appendix A summarizes the datasets used in Stage III.
@@
Table 1: Results on different academic task-oriented datasets. "Res.", "Trained" and "Data" represent the input image resolution, the number of trainable parameters, and the number of training instruction-response pairs, respectively. Both BLIP-2 (Li et al., 2023b) and InstructBLIP-2 (Dai et al., 2024) refer to the FlanT5XL (Chung et al., 2024) variants, while moondream2 refers to the 2024-04-02 model version. The symbol * denotes evaluations made with OCR tokens in the instruction, while refers to our own evaluation of the models, using the prompt "Provide a one-sentence caption for the provided image", or instead the prompt "Generate a highly detailed description for the provided image using multiple sentences" for the case of DOCCI (Onoe et al., 2024), without OCR tokens in the instruction. Pheye-xN denotes a Pheye model with dense cross-attention layers inserted every $N$ layers of the Phi 1.5 model. VQAv2 refers to the test-dev split, NoCaps (Agrawal et al., 2019) and TextVQA refer to the validation splits, and TextCaps and DOCCI refer to the test splits of the corresponding datasets. VQAv2 and TextVQA report VQA accuracy, while NoCaps, TextCaps, and DOCCI report CIDEr.
@@
We initially aimed for a model that can effectively describe images, including their finer details. We used ShareGPT4V-PT (Chen et al., 2023a), featuring 1,246K images with detailed descriptions, of which 570K are high resolution images from SAM (Kirillov et al., 2023). To reduce overfitting to a particular prompt, ten different captioning instructions were manually generated. One of these instructions was then randomly selected to be associated with each image-description pair.
@@
The second stage empowers Pheye to go beyond captioning, using a mixture of complex multi-modal instruction following examples: MIMIC-IT (Li et al., 2023a), LRV (Liu et al., 2023), SViT (Zhao et al., 2023) and LVIS (Wang et al., 2023). This comes to a total of 964K samples. For each sample with multiple instruction-response turns, a single turn was randomly selected.
@@
Previous stages used synthetic GPT-4 (Achiam et al., 2023) and GPT-4V (OpenAI, 2023) data for finetunning, which is naturally prone to noise and hallucinations. To alleviate this, we further finetunned our models on a mixture of academic task-oriented VQA and captioning datasets, as well as some synthetic multimodal instruction following examples, based on LLaVA 1.5-mix-665K. We specifically considered ST-VQA (Biten et al., 2019), TextVQA, and the LLaVAR finetunning dataset for better scene-text performance, removed the RefCOCO (Kazemzadeh et al., 2014) and VisualGenome (Krishna et al., 2017) datasets, and sampled 2 random turns from the VQAv2 (Antol et al., 2015) and GQA (Hudson and Manning, 2019) examples for each image. As a result, the model is only trained on $22.4\%$ of the VQAv2 and GQA examples in the original LLaVA 1.5 mix. We also sampled single turns from LLaVA, LLaVAR, OCR-VQA (Mishra et al., 2019) and OKVQA (Marino et al., 2019). From A-OKVQA (Schwenk et al., 2022), we randomly sampled unique image-question pairs.
@@ 3.2 Experimental Results
Table 1 summarizes our experimental results. With less training data and less trainable parameters, Pheye surpasses other generalist models of similar size in scence-text tasks, without using pre-extracted OCR tokens in the textual instructions. Note, for instance, that MobileVLM V2 (Chu et al., 2024) exhibits subpar performance on TextCaps, which suggests that the model relies heavily on the presence of OCR tokens in the instruction.
@@ 3.3 Assessing the Use of Fine Image Details
In order to further evaluate the impact that increasing the input image resolution has on tasks that involve the comprehension of fine-grained details, we followed a strategy similar to that of Zhang et al. (2023a) and partitioned the TextVQA validation set into three groups of approximately equal size, based on the relative size of the ground-truth bounding box $S=\frac{A_{bb}}{A_{total}}$, where $A_{bb}$ denotes the area of the answer bounding box, and $A_{total}$ denotes the total area of the image. Specifically, we divided the data into three tertiles: the bottom tertile (finer details), the middle tertile (medium), and the top tertile (broader entities). The selection of the ground-truth bounding box was based on the average string similarity with all the ground truth answers, using the longest contiguous matching subsequence algorithm.
@@ 4 Conclusions
Future work directions include investigating the use of different vision encoders, for instance incorporating strategies to process images at close to native resolutions and aspect ratios, e.g. building upon the approach proposed by Dehghani et al. (2024). We can also explore ways to increase the amount of training data, given that our method still requires training hundreds of millions of randomly initialized parameters. To address this, we could generate additional task-specific synthetic data, particularly for tasks that involve scene-text, where the amount of available human generated data is smaller. Building on the work of Zhang et al. (2023a), we could generate VQA examples that focus on finer image details, which have been shown to be crucial for performance.
@@
Another important limitation in the work reported on this paper concerns the fact that our experiments relied exclusively on English datasets. Multilingual models have shown potential in leveraging diverse datasets and providing more robust and versatile language understanding capabilities, which could be beneficial for creating VLMs that can handle a wider variety of tasks and languages. Future work can perhaps explore the use of efficient multilingual models like Qwen (Bai et al., 2023) to enhance our approach, although additional efforts would be required in the design of an effective mixture of multilingual data for training.
@@ Appendix B Preliminary Experiments Assessing Architectural Choices
The experiments followed the main principles of Flamingo (Alayrac et al., 2022), where the language and vision models were frozen to preserve their pre-training knowledge.
@@ B.1 Vision-Language Model Architectures
One initial experiment evaluated three different alternatives for combining the vision and language modalities, taking inspiration from the FROMAGe (Koh et al., 2023), LLaVA, and SmallCap (Ramos et al., 2023) neural architectures.
@@
Table 4: Comparison of different approaches for combining the vision and language modalities. The column "Train" denotes models in which the ViT was trained, and the column "Param." denotes the number of trainable parameters. "B@4" and "M" represent the BLEU-4 (Papineni et al., 2002) and METEOR (Banerjee and Lavie, 2005) metrics, respectively.
@@
To compare the effectiveness of the aforementioned three architectures, we trained and evaluated them on a captioning task using the COCO dataset (Lin et al., 2014). Specifically, we trained our models on the training split, using a cross-entropy loss, and evaluated them on the validation split. The ViT was initialized with pre-trained weights from a CLIP-ViT-B-32 model, while the language model was initialized with pre-trained weights from the GPT-2 base model (Radford et al., 2019). During training, the language model was kept frozen, while the connector module, comprising either a linear mapping in the first two architectures or cross-attention modules in the third, was trained. Additionally, we experimented with training the ViT. The models were trained for 5 epochs using the AdamW optimizer, with a batch size of 64 instances, and a learning rate of 1e-4 for the cross-attention modules and 1e-5 for the vision encoder. The GPT-2 model was trained to generate captions using the prefix "This image shows ", and the model was also prompted with this prefix during the evaluation stage.
@@
Similar findings to those reported in this section have also been reported by Laurençon et al. (2024), although these authors have also showed that LLaVA-style linear projections can outperform cross-attention if, besides the ViT, the language model is also carefully trained.
@@ B.2 Encoding High Resolution Images
Previous attempts at increasing the input resolution of VLMs mostly chose to finetune the vision encoder on higher resolution images (Chen et al., 2023b), which is costly due to the quadratic computational complexity of the attention mechanism, and which also requires large amounts of image-text data. To overcome this issue, we experimented with an approach similar to that of Li et al. (2023d), in which we scale up the input resolution by splitting the image into smaller patches that match the resolution of the vision encoder, afterwards encoding the smaller sub-images individually. To provide the model with global context, we also encode the full image as normal, concatenate all feature maps, and use the result as our image representation.
@@
The proposed method generates image representations with larger sequence lengths, which can be computationally demanding. To mitigate this issue, we explored the use of two resampler architectures, namely a 6-layer Perceiver Resampler, as introduced in Flamingo Alayrac et al. (2022), and the Monkey Resampler, which can be characterized as a single layer Perceiver Resampler that does not include a feed-forward module and residual connections for the query vectors. In our experiments, both modules used 257 query vectors with the same dimensionality of the vision encoder. Additionally, we compare the aforementioned resamplers, which encode images at 448x448 pixels, with an encoder that processes images at 224x224 pixels, and also with a 448x448 pixels encoder that does not compress the image features. For a fair comparison, the smaller 224x224 pixels encoder also includes one set of LoRA adapters.
Due to the fact that scaling up the resolution is likely to benefit tasks that involve scene-text and OCR the most, we pre-train our models on the LLaVAR (Zhang et al., 2023b) dataset for 1 epoch, finetune on the TextCaps (Sidorov et al., 2020) train split for 3 epochs, and finally evaluate on the TextCaps validation split. When finetuning our models, we use the prompt "This image shows ", similarly to the previous experiment. Our vision Transformer and language model are initialized from pre-trained CLIP-ViT-L-14 and GPT-2 large models, respectively, and are combined with cross-attention modules, following SmallCap. The use of larger a ViT and language model, when compared to the experiment described in the previous appendix, relates to the fact that tasks involving scene-text and OCR are more demanding, and hence slightly larger models can facilitate the assessment of differences between the architectural alternatives being compared. We freeze the ViT and language model, and only train the cross-attention modules, positional embeddings, LoRA, and resamplers, using a cross-entropy loss over non-prompt tokens. We use the AdamW optimizer, with a learning rate of 1e-4 for the pre-training stage and 2e-5 for the fine-tuning stage, and with a batch size of 64 in both stages. The results are shown in Table 5.
The experiment revealed that increasing the input resolution from 224x224 to 448x448 pixels increases captioning performance, independently from whether we compress the resulting visual tokens or not. There does not seem to be a large performance difference in how we compress the visual tokens, since the Monkey Resampler and the Perceiver Resampler have similar CIDEr (Vedantam et al., 2015) scores on the TextCaps validation split. However, this might not be the case for other tasks. Finally, the architecture with the best results is the one that increases the input resolution from 224x224 pixels to 448x448 pixels without compressing the vision features.

### v2

@@ 1 Introduction
The integration of visual capabilities as extensions to Large Language Models (LLMs) has led to the emergence of large Vision-Language Models (VLMs), currently excelling in tasks like image captioning and visual question answering. Notable examples include Flamingo (alayrac2022flamingo) or Monkey (li2023monkey), in this last case bringing improvements to tasks that involve scene-text by processing higher-resolution images. LLaVA-NeXT (liu2024llavanext) followed Monkey’s lead with similar enhancements, but at the cost of a quadratic increase in the computational complexity of the language model, as it skipped a resampler module that compresses large sequence lengths, from the encoder, into a fixed size of query vectors.
Recent focus has shifted towards smaller VLMs that can run on hardware-constrained devices. Models like MoE-LLaVA (lin2024moe) or moondream2 (moondream2) achieve impressive performance with a fraction of the parameters of their predecessors. This work further advances small VLMs with Pheye, i.e. a family of compact models that can process high-resolution images with fewer parameters and computational demands, expanding VLM applications to resource-limited environments where understanding fine details is crucial.
Specifically, Pheye employs a frozen instruction-tuned language model (li2023textbooks) in conjunction with a frozen pre-trained CLIP (radford2021learning) vision model, which are linked by dense cross-attention layers inserted before the language model’s layers. To process high-resolution images, we use two sets of LoRA (hu2021lora) adapters in the vision encoder, one for the global image and another for local high-resolution patches.
@@
Notably, Pheye is competitive with similarly sized models, particularly in tasks involving scene-text, such as TextVQA (singh2019towards). Pheye requires only a fraction of the training parameters, being more efficient in connecting the vision and language modalities, and in processing input images at higher resolutions. The source code associated to the experiments reported on this paper, as well as the trained models, are available from a public GitHub repository [^1] https://github.com/miguelscarv/pheye.
@@ 2.1 The Proposed Architecture
The vision encoder consists of a ViT with two sets of LoRA adapters, one for encoding global images and another for local high-resolution patches, as outlined in Appendix B.2. To better distinguish between global and local patch embeddings and improve convergence, we introduce two LayerNorm (ba2016layer) layers after the ViT, and before adding learned positional embeddings. These layers are applied separately, respectively processing the global and local patch embeddings.
@@
Given the computational efficiency, strong task performance, and reduced number of trainable parameters, we use cross-attention to combine modalities while keeping the language model frozen. Inspired by the Flamingo architecture (alayrac2022flamingo), we replace vanilla cross-attention modules with dense cross-attention modules, inserting them at regular intervals before the decoder layers, as shown in Figure 2. This design was motivated Flamingo’s demonstration that gated cross-attention outperforms vanilla cross-attention even when parameter counts are equal. However, we deviate from Flamingo’s gated cross-attention layers, as early experiments showed that the gating mechanism hindered convergence. To preserve the language model’s initial integrity, we initialize the output matrices of the dense cross-attention modules with values sampled from a normal distribution with a mean of zero and a variance approaching zero. This way, the cross-attention layers have a minor influence during the initial training epochs.
@@ 2.2 Analysis of the Computational Complexity
$\vskip-0.28436pt\mathrm{\mathbb{T}_{ViT}}=4ND^{2}+DN^{2}+8ND^{2}.\vskip-1.42271pt$ (1)
@@
We can compare the efficiency of our vision encoder against a ViT that operates at a resolution of $\times672$ pixels with a patch size of 14 pixels, resulting in $N=2305$ input tokens. Our method would process 10 images, where 1 is global and 9 are local sub-images, at a resolution of $\times224$ pixels with a patch size of 14, yielding $N^{\prime}=2570$ tokens. With $D=1280$, our method is approximately 1.02 as efficient as the alternative. While this corresponds to little to no improvement, our method does offer a significant advantage: it eliminates the need to fine-tune the underlying ViT, allowing us to focus solely on training the LoRA parameters to increase the input resolution.
@@ 3.1 Experimental Setup
Our experiments aimed to assess the effects of increasing the input image resolution, and to examine the impact of augmenting the frequency of dense cross-attention layers in the language model. To do this, we compare four model settings, varying the image input resolution between $\times448$ pixels and $\times672$ pixels, and adjusting the dense cross-attention interval to every 4 or 2 decoder layers.
In all four settings, we used a Phi 1.5 language model (li2023textbooks), finetunned on the SlimOrca (SlimOrca) instruction dataset. The ViT is initialized from a CLIP-ViT-H-14 model finetunned on DFN-5B (fang2023data).
The models were trained in separate stages, similarly to MoE-LLaVA (lin2024moe), using a cross-entropy loss over the output tokens. The different stages are described next, while Appendix A summarizes the datasets used in Stage III.
@@
Table 1: Results on different academic task-oriented datasets. "Res.", "Trained" and "Data" represent the input image resolution, the number of trainable parameters, and the number of training instruction-response pairs, respectively. Both BLIP-2 (li2023blip) and InstructBLIP-2 (dai2024instructblip) refer to the FlanT5XL (chung2024scaling) variants, while moondream2 refers to the 2024-04-02 model version. The symbol * denotes evaluations made with OCR tokens in the instruction, while refers to our own evaluation of the models, using the prompt "Provide a one-sentence caption for the provided image", or instead the prompt "Generate a highly detailed description for the provided image using multiple sentences" for the case of DOCCI (OnoeDocci2024), without OCR tokens in the instruction. Pheye-xN denotes a Pheye model with dense cross-attention layers inserted every NN layers of the Phi 1.5 model. VQAv2 refers to the test-dev split, NoCaps (agrawal2019nocaps) and TextVQA refer to the validation splits, and TextCaps and DOCCI refer to the test splits of the corresponding datasets. VQAv2 and TextVQA report VQA accuracy, while NoCaps, TextCaps, and DOCCI report CIDEr.
@@
We initially aimed for a model that can effectively describe images, including their finer details. We used ShareGPT4V-PT (chen2023sharegpt4v), featuring 1,246K images with detailed descriptions, of which 570K are high resolution images from SAM (kirillov2023segment). To reduce overfitting to a particular prompt, ten different captioning instructions were manually generated. One of these instructions was then randomly selected to be associated with each image-description pair.
@@
The second stage empowers Pheye to go beyond captioning, using a mixture of complex multi-modal instruction following examples: MIMIC-IT (li2023mimic), LRV (liu2023aligning), SViT (zhao2023svit) and LVIS (wang2023see). This comes to a total of 964K samples. For each sample with multiple instruction-response turns, a single turn was randomly selected.
@@
Previous stages used synthetic GPT-4 (achiam2023gpt) and GPT-4V (gpt4v) data for finetunning, which is naturally prone to noise and hallucinations. To alleviate this, we further finetunned our models on a mixture of academic task-oriented VQA and captioning datasets, as well as some synthetic multimodal instruction following examples, based on LLaVA 1.5-mix-665K. We specifically considered ST-VQA (biten2019scene), TextVQA, and the LLaVAR finetunning dataset for better scene-text performance, removed the RefCOCO (refcoco) and VisualGenome (krishna2017visual) datasets, and sampled 2 random turns from the VQAv2 (antol2015vqa) and GQA (hudson2019gqa) examples for each image. As a result, the model is only trained on $22.4\%$ of the VQAv2 and GQA examples in the original LLaVA 1.5 mix. We also sampled single turns from LLaVA, LLaVAR, OCR-VQA (ocrvqa) and OKVQA (marino2019ok). From A-OKVQA (schwenk2022okvqa), we randomly sampled unique image-question pairs.
@@ 3.2 Experimental Results
Table 1 summarizes our experimental results. With less training data and less trainable parameters, Pheye surpasses other generalist models of similar size in scence-text tasks, without using pre-extracted OCR tokens in the textual instructions. Note, for instance, that MobileVLM V2 (chu2024mobilevlm) exhibits subpar performance on TextCaps, which suggests that the model relies heavily on the presence of OCR tokens in the instruction.
@@ 3.3 Assessing the Use of Fine Image Details
In order to further evaluate the impact that increasing the input image resolution has on tasks that involve the comprehension of fine-grained details, we followed a strategy similar to that of zhang2023visual and partitioned the TextVQA validation set into three groups of approximately equal size, based on the relative size of the ground-truth bounding box $S=\frac{A_{bb}}{A_{total}}$, where $A_{bb}$ denotes the area of the answer bounding box, and $A_{total}$ denotes the total area of the image. Specifically, we divided the data into three tertiles: the bottom tertile (finer details), the middle tertile (medium), and the top tertile (broader entities). The selection of the ground-truth bounding box was based on the average string similarity with all the ground truth answers, using the longest contiguous matching subsequence algorithm.
@@ 4 Conclusions
Future work directions include investigating the use of different vision encoders, for instance incorporating strategies to process images at close to native resolutions and aspect ratios, e.g. building upon the approach proposed by dehghani2024patch. We can also explore ways to increase the amount of training data, given that our method still requires training hundreds of millions of randomly initialized parameters. To address this, we could generate additional task-specific synthetic data, particularly for tasks that involve scene-text, where the amount of available human generated data is smaller. Building on the work of zhang2023visual, we could generate VQA examples that focus on finer image details, which have been shown to be crucial for performance.
@@
Another important limitation in the work reported on this paper concerns the fact that our experiments relied exclusively on English datasets. Multilingual models have shown potential in leveraging diverse datasets and providing more robust and versatile language understanding capabilities, which could be beneficial for creating VLMs that can handle a wider variety of tasks and languages. Future work can perhaps explore the use of efficient multilingual models like Qwen (qwen) to enhance our approach, although additional efforts would be required in the design of an effective mixture of multilingual data for training.
@@ Appendix B Preliminary Experiments Assessing Architectural Choices
The experiments followed the main principles of Flamingo (alayrac2022flamingo), where the language and vision models were frozen to preserve their pre-training knowledge.
@@ B.1 Vision-Language Model Architectures
One initial experiment evaluated three different alternatives for combining the vision and language modalities, taking inspiration from the FROMAGe (koh2023grounding), LLaVA, and SmallCap (ramos2023smallcap) neural architectures.
@@
Table 4: Comparison of different approaches for combining the vision and language modalities. The column "Train" denotes models in which the ViT was trained, and the column "Param." denotes the number of trainable parameters. "B@4" and "M" represent the BLEU-4 (papineni-etal-2002-bleu) and METEOR (banerjee-lavie-2005-meteor) metrics, respectively.
@@
To compare the effectiveness of the aforementioned three architectures, we trained and evaluated them on a captioning task using the COCO dataset (lin2014microsoft). Specifically, we trained our models on the training split, using a cross-entropy loss, and evaluated them on the validation split. The ViT was initialized with pre-trained weights from a CLIP-ViT-B-32 model, while the language model was initialized with pre-trained weights from the GPT-2 base model (radford2019language). During training, the language model was kept frozen, while the connector module, comprising either a linear mapping in the first two architectures or cross-attention modules in the third, was trained. Additionally, we experimented with training the ViT. The models were trained for 5 epochs using the AdamW optimizer, with a batch size of 64 instances, and a learning rate of 1e-4 for the cross-attention modules and 1e-5 for the vision encoder. The GPT-2 model was trained to generate captions using the prefix "This image shows ", and the model was also prompted with this prefix during the evaluation stage.
@@
Similar findings to those reported in this section have also been reported by laurenccon2024matters, although these authors have also showed that LLaVA-style linear projections can outperform cross-attention if, besides the ViT, the language model is also carefully trained.
@@ B.2 Encoding High Resolution Images
Previous attempts at increasing the input resolution of VLMs mostly chose to finetune the vision encoder on higher resolution images (chen2023pali), which is costly due to the quadratic computational complexity of the attention mechanism, and which also requires large amounts of image-text data. To overcome this issue, we experimented with an approach similar to that of li2023monkey, in which we scale up the input resolution by splitting the image into smaller patches that match the resolution of the vision encoder, afterwards encoding the smaller sub-images individually. To provide the model with global context, we also encode the full image as normal, concatenate all feature maps, and use the result as our image representation.
@@
The proposed method generates image representations with larger sequence lengths, which can be computationally demanding. To mitigate this issue, we explored the use of two resampler architectures, namely a 6-layer Perceiver Resampler, as introduced in Flamingo alayrac2022flamingo, and the Monkey Resampler, which can be characterized as a single layer Perceiver Resampler that does not include a feed-forward module and residual connections for the query vectors. In our experiments, both modules used 257 query vectors with the same dimensionality of the vision encoder. Additionally, we compare the aforementioned resamplers, which encode images at 448x448 pixels, with an encoder that processes images at 224x224 pixels, and also with a 448x448 pixels encoder that does not compress the image features. For a fair comparison, the smaller 224x224 pixels encoder also includes one set of LoRA adapters.
Due to the fact that scaling up the resolution is likely to benefit tasks that involve scene-text and OCR the most, we pre-train our models on the LLaVAR (zhang2023llavar) dataset for 1 epoch, finetune on the TextCaps (sidorov2020textcaps) train split for 3 epochs, and finally evaluate on the TextCaps validation split. When finetuning our models, we use the prompt "This image shows ", similarly to the previous experiment. Our vision Transformer and language model are initialized from pre-trained CLIP-ViT-L-14 and GPT-2 large models, respectively, and are combined with cross-attention modules, following SmallCap. The use of larger a ViT and language model, when compared to the experiment described in the previous appendix, relates to the fact that tasks involving scene-text and OCR are more demanding, and hence slightly larger models can facilitate the assessment of differences between the architectural alternatives being compared. We freeze the ViT and language model, and only train the cross-attention modules, positional embeddings, LoRA, and resamplers, using a cross-entropy loss over non-prompt tokens. We use the AdamW optimizer, with a learning rate of 1e-4 for the pre-training stage and 2e-5 for the fine-tuning stage, and with a batch size of 64 in both stages. The results are shown in Table 5.
The experiment revealed that increasing the input resolution from 224x224 to 448x448 pixels increases captioning performance, independently from whether we compress the resulting visual tokens or not. There does not seem to be a large performance difference in how we compress the visual tokens, since the Monkey Resampler and the Perceiver Resampler have similar CIDEr (vedantam2015cider) scores on the TextCaps validation split. However, this might not be the case for other tasks. Finally, the architecture with the best results is the one that increases the input resolution from 224x224 pixels to 448x448 pixels without compressing the vision features.